The actual selling expenses incurred in March 2020 by Fallon Company are as follows.

Variable Expenses

Fixed Expenses

Sales commissions $11,272 Sales salaries $34,900
Advertising 10,552 Depreciation 7,500
Travel 8,685 Insurance 1,600
Delivery 3,534


(a) Prepare a flexible budget performance report for March, assuming that March sales were $173,700. Variable costs and their percentage relationship to sales are sales commissions 6%, advertising 6%, traveling 5%, and delivery 2%. Fixed selling expenses will consist of sales salaries $34,900, Depreciation on delivery equipment $7,500, and insurance on delivery equipment $1,600. (List variable costs before fixed costs.)

(b) Prepare a flexible budget performance report, assuming that March sales were $181,000. (List variable costs before fixed costs.)
Answer #1

Answer-a)-

FALLON COMPANY
Selling Expense Flexible Budget Report
For the month ended March 31,2017
Particulars Flexible Budget Actual Costs Differences Remark
$ $ $
Variable Expenses:-
Sales commissions 6% $173700*6%=$10422 11272 -850 Unfavorable
Advertising 6% $173700*6%=$10422 10552 -130 Unfavorable
Travel 5% $173700*5%=$8685 8685 0 None
Delivery 2% $173700*2%=$3474 3534 -60 Unfavorable
Total Variable expenses (a) 33003 34043 -1040 Unfavorable
Fixed Expenses:-
Sales salaries 34900 34900 0 None
Depreciation 7500 7500 0 None
Insurance 1600 1600 0 None
Total Fixed expenses (b) 44000 44000 0 None
Total Expenses (c=a+b) 77003 78043 -1040 Unfavorable

b)-

FALLON COMPANY
Selling Expense Flexible Budget Report
For the month ended March 31,2017
Particulars Flexible Budget Actual Costs Differences Remark
$ $ $
Variable Expesnes:-
Sales commissions 6% $181000*6%=$10860 11272 -412 Unfavorable
Advertising 6% $181000*6%=$10860 10552 308 Favorable
Travel 5% $181000*5%=$9050 8685 365 Favorable
Delivery 2% $181000*2%=$3620 3534 86 Favorable
Total Variable expenses (a) 34390 34043 347 Favorable
Fixed Expenses:-
Sales salaries 34900 34900 0 None
Depreciation 7500 7500 0 None
Insurance 1600 1600 0 None
Total Fixed expenses (b) 44000 44000 0 None
Total Expenses (c=a+b) 78390 78043 347 Favorable